OpenStudy (anonymous):

HELP PLEASE! T varies inversely with the cube of W When W is 3 T is 1/9 find the value of W when T is 1/243

OpenStudy (anonymous):

@LarsEighner @ikram002p

OpenStudy (ikram002p):

post the whol Qn plz :)

OpenStudy (anonymous):

@ikram002p here you go

OpenStudy (ikram002p):

\(T \text{ }\alpha \text{ } \dfrac{1}{w}\) means \(T \text{ }= \text{ } \dfrac{k}{w}\) s.t k is constant \( \dfrac{1}{9} \text{ }= \text{ } \dfrac{k}{3}\) so \(k=\dfrac{1}{3}\) equation become \(T = \dfrac{1}{3w}\)

OpenStudy (ikram002p):

so \(T = \dfrac{1}{3w}\) \(\dfrac{1}{243} = \dfrac{1}{3w}\) solve for w

OpenStudy (ikram002p):

i got 81